# Partition function zeros of quantum many-body systems

### v2

@@ Abstract
We present a method for calculating the Yang-Lee partition function zeros of a translationally invariant model of lattice fermions, exemplified by the Hubbard model. The method rests on a theorem involving the single electron self-energy $\Sigma_{\sigma}(\vec{k},i\omega_{n})$ in the imaginary time Matsubara formulation. The theorem maps the Yang-Lee zeros to a set of wavevector and spin labeled virtual energies $\xi_{\vec{k}{\sigma}}$. These, thermodynamically derived virtual energies, are solutions of equations involving the self-energy at corresponding $\vec{k}{\sigma}$’s. Examples of the method in simplified situations are provided.
@@ 1 Introduction
Our interest is in the study of quantum many-body systems, such as the Hubbard model. Locating the zeros and extracting meaningful information for these systems has proceeded at a sedate pace. Here the major difficulty is the absence of systematic analytical methods. For numerical studies, a combination of the very high precision required for locating the zeros and the exponential growth of the Hilbert spaces with the system size, hamper direct enumeration techniques. Quantum Monte Carlo (QMC) methods for studying the repulsive Hubbard model with a complex chemical potential have been used earlier [9, 10, 11, 12]. However the full spectrum of zeros seems difficult to obtain from QMC. There have also been recent studies of partition function zeros for related quantum many-body models in a variety of contexts, these include decoherence studies[13], quantum quenches[14] and in quantum computation [15]. These indicate a renewal of interest in this important subject.
The main objective of this paper is to introduce a new analytical method for locating the zeros of the grand partition function of the Hubbard model (see Eq. (5))- in any dimension. This method requires a knowledge of the self energy in the Matsubara imaginary time framework[16]. Perturbative knowledge of the self-energy may be used to give approximate location of the zeros. The method generalizes to nearby models in a straightforward way. The a few simple examples are provided below.
@@ 1.2 Plan of the paper
In Section(2) we provide basic definitions of the Hamiltonian, the grand partition function and the fugacity. We define the virtual energies $\xi_{\vec{k}{\sigma}}$, and note some properties of these obtained from general considerations. In Section(3) we define the Matsubara finite temperature Greens functions, their un-normalized versions and the self-energy. In Section(4) the main theorem of this work is proposed and established. The equation determining the virtual energy in terms of the self-energy is recorded in Eq. (28). In Section(5) we illustrate the application of the theorem to the problem of an exactly solvable single site Hubbard model. Here the role of lifting of degeneracy of the single site energy is also studied in detail. The exact self energy is expanded to ${\cal O}(U^{2})$, and to this order the perturbative results for $\xi$ are shown to agree with the exact ones. This somewhat long section is mainly addressed to readers interested in details, and may be skipped by others. In Section(6) we use the first order (Hartree) approximation to the self energy and extract the virtual energies $\xi_{\vec{k}{\sigma}}$ to ${\cal O}(U)$ for both signs of the interaction for the 1-d Hubbard ring of length 6. This is familiar as the idealized Benzene molecule. These virtual zeros are compared with exact numerical values obtained by computing the grand partition function exactly using symbolic methods. In Section(7) we note a few conclusions and prospects for further work.
In Appendix(A) we recall the well known results of Luttinger and Ward[20] for the grand partition function. Their result expresses the result of an all-orders summation of perturbation theory. We observe that the key function $\Xi(\mu)$ in Eq. (27)- whose zeros are all the virtual energies- is a multiplicative factor of the Luttinger-Ward partition function.
@@ 2 Basic Hamiltonian and the partition function
where $E_{n}$ and $N_{n}$ are the eigenvalues of $H$ and particle number for the state $n$ (detailed definitions are given after Eq. (29)), $Z_{N}$ is the canonical partition function for $N$ particles and $z$ is the fugacity
$z(\mu)=e^{\beta\mu}.$ (7)
$Z(\mu)$ is thus a polynomial in the fugacity $z$ of degree $2N_{s}$, and by the fundamental theorem of algebra, it has exactly $2N_{s}$ zeros in the complex $z$ plane- and hence expressible as $Z(\mu)=c_{0}\prod_{\alpha=1}^{2N_{s}}(z-z_{\alpha})$, where $z_{\alpha}$ are the Yang-Lee zeros of the given model. We will find it more convenient to work in the complex $\mu$ plane rather than with the fugacity $z$. We note the periodicity $z(\mu)=z(\mu+i\frac{2\pi}{\beta})$ and therefore the complex $\mu$ plane can be decomposed into infinite equivalent strips of width $i\frac{2\pi}{\beta}$. We restrict our attention to a conveniently chosen strip
@@ 2.1 Definition of the virtual energies $\xi_{\vec{k}{\sigma}}$
$Z(\mu)=\prod_{\vec{k}{\sigma}}(1+\frac{e^{\beta\mu}}{e^{\beta(\xi_{\vec{k}{\sigma}}+\frac{U}{2})}}).$ (9)
@@
$\mu$$\to$ $-i\frac{\pi}{\beta}+\frac{U}{2}+\xi_{\vec{k}{\sigma}},\mbox{ or equivalently }$
$z_{\alpha}$$\to$ $-e^{\beta(\xi_{\vec{k}{\sigma}}+\frac{U}{2})}$ (10)
where Eq. (10) anticipates a mapping $\alpha\leftrightarrow\vec{k}{\sigma}$, between the unlabeled Yang-Lee zeros $z_{\alpha}$ and the labeled virtual energies $\xi$. We remark that while Eq. (9) has the same form as that of an ideal Fermi gas with energies $\xi_{\vec{k}{\sigma}}+\frac{U}{2}$, the latter are non-trivial complex functions of $T,U$. For an interacting system $\xi_{\vec{k}{\sigma}}$ are not viewed as eigenvalues of any Hermitean Hamiltonian, and are allowed to be non-real. For this reason these may be regarded as thermodynamically derived virtual energies.
@@
The convention used in defining the strip ${\cal S}$ in Eq. (8), and the range of the imaginary part of $\xi$ in Eq. (11), are such that complex conjugate roots in the fugacity variable $z$ Eq. (7) lead to complex conjugate energies $\xi$. The partition function being a polynomial in $z$ with real coefficients, the roots in $z$ are either real or come in complex conjugate pairs. It follows that the virtual energies are either real or they occur in complex conjugate pairs, i.e. if $\xi$ is a virtual energy, then so is $\xi^{*}$.
@@
  • A second sumrule can be readily found by comparing the results of the single particle sector $N=1$, i.e the coefficient of $e^{\beta\mu}$ in Eq. (9). Its exact value is found using the observation that the two-particle interaction $U$ is inoperative in this sector. We find the second sum-rule:
    $\sum_{\vec{k}{\sigma}}e^{-\beta(\xi_{\vec{k}{\sigma}}+\frac{U}{2})}=\sum_{\vec{k}{\sigma}}e^{-\beta\varepsilon_{\vec{k}{\sigma}}}.$ (15)
@@
3 The Greens function and Self energy in Matsubara frequencies
@@
Defining the Dyson self energy $\Sigma$ as
@@
It should be noted that in $\Sigma_{\sigma}(\vec{k},i\omega_{r}|\mu)$, every occurrence of $i\omega_{r}$ is in the combination $i\omega_{r}+\mu$. In terms with $r\neq 0$ we note that $\omega_{r}-\omega_{0}=r\frac{2\pi}{\beta}$, this shift can be absorbed into $\mu$ and using the periodicity $Z(\mu)=Z(\mu+ir\frac{2\pi}{\beta})$. We also note that Eq. (302122) implies the periodicity in the complex $\mu$ plane for integer $s$
@@
  • In principle the self-energy $\Sigma_{\sigma}(\vec{k},i\omega_{r}|\mu)$ can be calculated to arbitrary orders in the interaction, using the standard rules of the Feynman-Dyson-Matsubara perturbation theory in the imaginary time[16, 17, 20]. Given an approximation for $\Sigma$ to a certain order in the coupling, the theorem can be used to find the corresponding (unlabeled) Yang-Lee zero of the partition function, or more precisely, the momentum-labeled virtual energy $\xi_{\vec{k}{\sigma}}$.
@@
    $\xi_{\vec{k}{\sigma}}=\varepsilon_{\vec{k}{\sigma}}-\frac{U}{2}+\Sigma_{{\sigma}}(\vec{k},i\frac{\pi}{\beta}|\xi_{k{\sigma}}+\frac{U}{2}-i\frac{\pi}{\beta}).$ (28)
@@ 5.2 Exact Green’s function in the Atomic Limit
$\xi_{i{\sigma}}=\frac{U}{2}+\varepsilon_{i0}+\frac{1}{\beta}\log\left(\cosh\beta\delta_{\varepsilon}+{\sigma}\{\sinh^{2}(\beta\delta_{\varepsilon})+1-e^{-\beta U}\}^{\frac{1}{2}}\right)$ (42)
@@
which can be rewritten in the form of a single term with a self energy
@@
It is easily seen that $\bar{G}_{\sigma}(i\omega_{0})$ is a holomorphic function of $\mu$ in the strip ${\cal S}$ Eq. (8) due to the vanishing of the numerator at every location of the vanishing denominators.
@@ 5.2.1 Roots for small U: Non degenerate case $\delta_{\varepsilon}>0$
$\xi_{\uparrow}=\epsilon_{0}+\delta_{\varepsilon}+\frac{U}{2}{\coth\beta\delta_{\varepsilon}}-\beta U^{2}\frac{1}{8}\frac{\cosh\beta\delta_{\varepsilon}}{\sinh^{3}\beta\delta_{\varepsilon}}+{\cal O}(U^{3})$ (47)
@@ 5.2.3 Perturbative Roots from $G$-The non-degenerate case
We next show how the small U roots given in Eq. (4748) can be obtained using the perturbative expansion of the Dyson self energy, as remarked below the theorem, i.e. the condition $\Psi_{\vec{k}{\sigma}}(\mu)=0$. Recall that
@@
For the purpose of this calculation we use the convenient variable $x$ which is related to the virtual energy through $\xi=x-\frac{U}{2}$. We next expand the exact occupation $n_{\sigma}$ appearing above, using Eq. (43), in a series in $U$. In terms of the Fermi function
@@ 5.2.4 Degenerate case
$x_{\pm}=\varepsilon_{0}\pm\sqrt{\frac{U}{\beta}}+\frac{U}{2}+{\cal O}(U^{3/2}),$ (65)
@@
6 Virtual energies from self energy to ${\cal O}(U)$.
The self energy of the Hubbard model on $N_{s}$ sites to ${\cal O}(U)$ is readily obtained from standard perturbation theory [17] as
@@
$\xi_{\vec{k}{\sigma}}$ $=$ $\varepsilon_{\vec{k}{\sigma}}-\frac{U}{2}+\frac{U}{N_{s}}\sum_{p}\frac{1}{1-e^{\beta(\varepsilon_{p{\bar{\sigma}}}-\xi_{k{\sigma}}-\frac{U}{2})}}+{\cal O}(U^{2}).$ (67)
@@
where we dropped a term of the ${\cal O}(U^{2})$ arising from expanding the $\coth$. The dropped term can be added back into other second order terms in the self energy. From the structure of Eq. (68) and since $\sum_{p}\varepsilon_{p{\bar{\sigma}}}=0$, we see that the sum-rule Eq. (14) is automatically satisfied.
@@
Figure 3: Comparison between the exact virtual energies found from the partition function zeros (red dots) and the solutions of Eq. (68) (blue dots) at four different repulsive values of $U$ marked in the plots.
@@
The formation of bound (Cooper) pairs in this case is similar to that of electron-electron binding posited in Dyson’s well known argument[19] about the instability of quantum electrodyamics with a flipped sign of $e^{2}$. The resulting non-convergence of perturbation theory in field theory, is based on the instability of the ground state due to pair formation, under the assumption of an infinite system. For the conventionally measured variables (energy, susceptibilities etc) related effects can only be seen in the limit of a large number of particles- i.e. in the thermodynamic limit. In the present case, we see that the distinct behaviour of the virtual states between the repulsive and attractive cases, namely the energies going from real values for repulsion to complex pairs for attraction. This striking change is visible even for small systems.
Figure 4: Attractive interactions All the attractive cases lead to complex virtual energies, as expected from arguments given in the text. Complex plots of the virtual energies with the real (imaginary) part on the x (y) axis. We compare the exact virtual energies found from the partition function zeros (red dots) and the solutions of Eq. (68) (blue dots) at four different attractive values of $U$ marked in the plots. We picture the degenerate roots by introducing a slight displacement in the real part. The formation of bound (Cooper type) pairs for $U<0$ shrink the bandwidth of the real part, these plots show such a tendency already at $U$=-$0.5$.

### v3

@@ Abstract
We present a new method for calculating the Yang-Lee partition function zeros of a translationally invariant model of lattice fermions, exemplified by the Hubbard model. The method rests on a theorem involving the usual single electron self-energy $\Sigma_{\sigma}(\vec{k},i\omega_{n}|\mu)$ with chemical potential $\mu$, in the imaginary time Matsubara formulation. The theorem maps the Yang-Lee zeros to a set of wavevector and spin labeled virtual energies $\xi_{\vec{k}{\sigma}}$. These, thermodynamically derived virtual energies, are solutions of the set of equations $\xi_{k{\sigma}}=\varepsilon_{\sigma}({\vec{k}})-\frac{1}{2}U+\Sigma_{\sigma}(\vec{k},i\frac{\pi}{\beta}|\xi_{k{\sigma}}+\frac{1}{2}U-i\frac{\pi}{\beta})=0$. Examples of the method in simplified situations are provided.
@@ 1 Introduction
Our interest is in the study of quantum many-body systems, such as the Hubbard model (see Eq. (5)). Locating the zeros and extracting meaningful information for these physically important systems has proceeded at a sedate pace. Here the major difficulty is the absence of systematic analytical methods. For numerical studies, a combination of the very high precision required for locating the zeros and the exponential growth of the Hilbert spaces with the system size, hamper direct enumeration techniques. Quantum Monte Carlo (QMC) methods for studying the repulsive Hubbard model with a complex chemical potential have been used earlier [9, 10, 11, 12]. However the full spectrum of zeros seems difficult to obtain from QMC. Studies of related problems employing tensor networks show promise in overcoming some of these limitations [13, 14]. There have also been recent studies of partition function zeros for related quantum many-body models in a variety of contexts, these include decoherence studies[15], quantum quenches[16] and in quantum computation [17]. These indicate a renewal of interest in this important subject.
The main objective of this paper is to introduce a new analytical method for locating the zeros of the grand partition function of the Hubbard model (see Eq. (5))- in any dimension. This method requires a knowledge of the self-energy in the Matsubara imaginary time framework[18]. While the self-energy is a highly non-trivial object to compute, a perturbative knowledge of the self-energy is available and may be used to give approximate location of the zeros. The method generalizes to nearby models in a straightforward way. The a few simple examples are provided below.
@@ 1.2 Plan of the paper
In Section(2) we provide basic definitions of the Hamiltonian, the grand partition function and the fugacity. We define the virtual energies $\xi_{\vec{k}{\sigma}}$, and note some properties of these obtained from general considerations. In Section(3) we define the Matsubara finite temperature Greens functions, their un-normalized versions and the self-energy. In Section(4) the main theorem of this work is proposed and established. The equation determining the virtual energy in terms of the self-energy is recorded in Eq. (29). In Section(5) we illustrate the application of the theorem to the problem of an exactly solvable single site Hubbard model. Here the role of lifting of degeneracy of the single site energy is also studied in detail. The exact self-energy is expanded to ${\cal O}(U^{2})$, and to this order the perturbative results for $\xi$ are shown to agree with the exact ones. This somewhat long section is mainly addressed to readers interested in details, and may be skipped by others. In Section(6) we use the first order (Hartree) approximation to the self-energy and extract the virtual energies $\xi_{\vec{k}{\sigma}}$ to ${\cal O}(U)$ for both signs of the interaction for the 1-d Hubbard ring of length 6. This is familiar as the idealized Benzene molecule. These virtual zeros are compared with exact numerical values obtained by computing the grand partition function exactly using symbolic methods. In Section(7) we note a few conclusions and prospects for further work.
In Appendix(A) we recall the well known results of Luttinger and Ward[23] for the grand partition function. Their result neatly expresses the result of an all-orders summation of perturbation theory. We observe that the key function $\Xi(\mu)$ in Eq. (28)- whose zeros are all the virtual energies- is a multiplicative factor of the Luttinger-Ward partition function.
@@ 2 Basic Hamiltonian and the partition function
where $E_{n}$ and $N_{n}$ are the eigenvalues of $H$ and particle number for the state $n$ (detailed definitions are given after Eq. (30)), $Z_{N}$ is the canonical partition function for $N$ particles and $z$ is the fugacity $z(\mu)=e^{\beta\mu}.$ $Z(\mu)$ is thus a polynomial in the fugacity $z$ of degree $2N_{s}$, and by the fundamental theorem of algebra, it has exactly $2N_{s}$ zeros in the complex $z$ plane- and hence expressible as $Z(\mu)=c_{0}\prod_{\alpha=1}^{2N_{s}}(z-z_{\alpha})$, where $z_{\alpha}$ are the Yang-Lee zeros of the given model. We will find it more convenient to work in the complex $\mu$ plane rather than with the fugacity $z$. We note the periodicity $z(\mu)=z(\mu+i\frac{2\pi}{\beta})$ and therefore the complex $\mu$ plane can be decomposed into infinite equivalent strips of width $i\frac{2\pi}{\beta}$. We restrict our attention to a conveniently chosen strip
@@ 2.1 Definition of the virtual energies $\xi_{\vec{k}{\sigma}}$
$Z(\mu)=\prod_{\vec{k}{\sigma}}(1+\frac{e^{\beta\mu}}{e^{\beta(\xi_{\vec{k}{\sigma}}+\frac{1}{2}U)}}).$ (8)
@@
$\mu$$\to$ $-i\frac{\pi}{\beta}+\frac{1}{2}U+\xi_{\vec{k}{\sigma}},\mbox{ or equivalently }$
$z_{\alpha}$$\to$ $-e^{\beta(\xi_{\vec{k}{\sigma}}+\frac{1}{2}U)}$ (9)
where Eq. (9) anticipates a mapping $\alpha\leftrightarrow\vec{k}{\sigma}$, between the unlabeled Yang-Lee zeros $z_{\alpha}$ and the labeled virtual energies $\xi$. We remark that while Eq. (8) has the same form as that of an ideal Fermi gas with energies $\xi_{\vec{k}{\sigma}}+\frac{1}{2}U$, the latter are non-trivial complex functions of $T,U$. For an interacting system $\xi_{\vec{k}{\sigma}}$ are not viewed as eigenvalues of any Hermitean Hamiltonian, and are allowed to be non-real. For this reason these may be regarded as thermodynamically derived virtual energies.
@@
The convention used in defining the strip ${\cal S}$ in Eq. (7), and the range of the imaginary part of $\xi$ in Eq. (10), are such that complex conjugate roots in the fugacity variable $z=e^{\beta\mu}$ lead to complex conjugate energies $\xi$. The partition function being a polynomial in $z$ with real coefficients, the roots in $z$ are either real or come in complex conjugate pairs. It follows that the virtual energies are either real or they occur in complex conjugate pairs, i.e. if $\xi$ is a virtual energy, then so is $\xi^{*}$.
@@
  • A second sum-rule can be readily found by comparing the results of the single particle sector $N=1$, i.e the coefficient of $e^{\beta\mu}$ in Eq. (8). Its exact value is found using the observation that the two-particle interaction $U$ is inoperative in this sector. We find the second sum-rule:
    $e^{-\beta\frac{1}{2}U}\sum_{\vec{k}{\sigma}}e^{-\beta\xi_{\vec{k}{\sigma}}}=\sum_{\vec{k}{\sigma}}e^{-\beta\varepsilon_{\vec{k}{\sigma}}}.$ (14)
    We now define $U_{eff}$ as a function of the parameters $U,\beta$ and also the computed $\{\xi\}$. It is given by
    $U_{eff}(U,\beta,\{\xi\})=\frac{2}{\beta}\log\left(\frac{\sum_{k{\sigma}}e^{-\beta\xi_{k{\sigma}}}}{\sum_{k{\sigma}}e^{-\beta\varepsilon_{k{\sigma}}}}\right).$ (15)
    The sum-rule Eq. (14) may be cast in the form
    $U_{eff}(U,\beta,\{\xi\})=U.$ (16)
    Clearly if the $\xi$’s are exactly found, then the sum-rule is satisfied exactly. If $\xi$’s are only found approximately, then the quality of the approximation is reflected in the difference between $U_{eff}$ and $U$. This idea is illustrated below in Fig. (34).
@@
3 The Greens function and Self-energy in Matsubara frequencies
@@
Defining the Dyson self-energy $\Sigma$ as
@@
It should be noted that in $\Sigma_{\sigma}(\vec{k},i\omega_{r}|\mu)$, every occurrence of $i\omega_{r}$ is in the combination $i\omega_{r}+\mu$. In terms with $r\neq 0$ we note that $\omega_{r}-\omega_{0}=r\frac{2\pi}{\beta}$, this shift can be absorbed into $\mu$ and using the periodicity $Z(\mu)=Z(\mu+ir\frac{2\pi}{\beta})$. We also note that Eq. (312223) implies the periodicity in the complex $\mu$ plane for integer ‘$s$
@@
  • In principle the self-energy $\Sigma_{\sigma}(\vec{k},i\omega_{r}|\mu)$ can be calculated to arbitrary orders in the interaction, using the standard rules of the Feynman-Dyson-Matsubara perturbation theory in the imaginary time[18, 19, 23]. Given an approximation for $\Sigma$ to a certain order in the coupling, the theorem can be used to find the corresponding momentum-labeled virtual energy $\xi_{\vec{k}{\sigma}}$, and from their total set, the (unlabeled) Yang-Lee zeros of the partition function.
@@
    $\xi_{\vec{k}{\sigma}}=\varepsilon_{\vec{k}{\sigma}}-\frac{1}{2}U+\Sigma_{{\sigma}}(\vec{k},i\frac{\pi}{\beta}|\xi_{k{\sigma}}+\frac{1}{2}U-i\frac{\pi}{\beta}).$ (29)
@@ 5.2 Exact Green’s function in the Atomic Limit
$\xi_{i{\sigma}}=\frac{1}{2}U+\varepsilon_{i0}+\frac{1}{\beta}\log\left(\cosh\beta\delta_{\varepsilon}+{\sigma}\{\sinh^{2}(\beta\delta_{\varepsilon})+1-e^{-\beta U}\}^{\frac{1}{2}}\right)$ (43)
@@
which can be rewritten in the form of a single term with a self-energy
@@
It is easily seen that $\bar{G}_{\sigma}(i\omega_{0})$ is a holomorphic function of $\mu$ in the strip ${\cal S}$ Eq. (7), due to the vanishing of the numerator at every location of the vanishing denominators.
@@ 5.2.1 Roots for small U: Non degenerate case $\delta_{\varepsilon}>0$
$\xi_{\uparrow}=\epsilon_{0}+\delta_{\varepsilon}+\frac{1}{2}U{\coth\beta\delta_{\varepsilon}}-\beta U^{2}\frac{1}{8}\frac{\cosh\beta\delta_{\varepsilon}}{\sinh^{3}\beta\delta_{\varepsilon}}+{\cal O}(U^{3})$ (48)
@@ 5.2.3 Perturbative Roots from $G$-The non-degenerate case
We next show how the small U roots given in Eq. (4849) can be obtained using the perturbative expansion of the Dyson self-energy, as remarked below the theorem, i.e. the condition $\Psi_{\vec{k}{\sigma}}(\mu)=0$. Recall that
@@
For the purpose of this calculation we use the convenient variable $x$ which is related to the virtual energy through $\xi=x-\frac{1}{2}U$. We next expand the exact occupation $n_{\sigma}$ appearing above, using Eq. (44), in a series in $U$. In terms of the Fermi function
@@ 5.2.4 Degenerate case
$x_{\pm}=\varepsilon_{0}\pm\sqrt{\frac{U}{\beta}}+\frac{1}{2}U+{\cal O}(U^{3/2}),$ (66)
@@
6 Virtual energies from self-energy to ${\cal O}(U)$.
The self-energy of the Hubbard model on $N_{s}$ sites to ${\cal O}(U)$ is readily obtained from standard perturbation theory [19] as
@@
$\xi_{\vec{k}{\sigma}}$ $=$ $\varepsilon_{\vec{k}{\sigma}}-\frac{1}{2}U+\frac{U}{N_{s}}\sum_{p}\frac{1}{1-e^{\beta(\varepsilon_{p{\bar{\sigma}}}-\xi_{k{\sigma}}-\frac{1}{2}U)}}+{\cal O}(U^{2}).$ (68)
@@
where we dropped a term of the ${\cal O}(U^{2})$ arising from expanding the $\coth$. The dropped term is added back to other second order terms in the self-energy, and the results of such a calculation to ${\cal O}(U^{2})$ will be reported later [20]. From the structure of Eq. (69) and since $\sum_{p}\varepsilon_{p{\bar{\sigma}}}=0$, we see that the sum-rule Eq. (13) is automatically satisfied.
@@
Figure 3: Comparison between the exact virtual energies found from the partition function zeros (red dots) and the solutions of the first order equations Eq. (69) (blue dots) at four different repulsive values of $U$ marked in the plots. The $x$ axis label index “$\nu$” is as defined in Fig. (2) for the Hartree solution and also the sorting index for the energies found from ED. For the displayed $U$ and $\beta$, the solutions of Eq. (69) can be used to compute $U_{eff}$ (see Eq. (1516)). The pairs $(U,U_{eff})$ are $(0.25,0.149),(0.5,0.289),(1.0,0.55),(1.5,0.804)$, indicating that the first order approximation worsens with increasing $U$.
@@
The formation of bound (Cooper) pairs in this case is similar to that of electron-electron binding posited in Dyson’s well known argument[22] about the instability of quantum electrodyamics with a flipped sign of $e^{2}$. The resulting non-convergence of perturbation theory in field theory, is based on the instability of the ground state due to pair formation, under the standard assumption of an infinite system. For the conventionally measured variables (energy, susceptibilities etc) related effects can only be seen in the limit of a large number of particles- i.e. in the thermodynamic limit. In the present case, we see that the distinct behaviour of the virtual states between the repulsive and attractive cases, namely the energies going from real values for repulsion to complex pairs for attraction. This striking change is visible even for small systems.
[…]